# F3D-Gaus: Feed-forward 3D-aware Generation on ImageNet with Cycle-Consistent Gaussian Splatting

### v1

@@ Abstract
This paper tackles the problem of generalizable 3D-aware generation from monocular datasets, e.g., ImageNet [10]. The key challenge of this task is learning a robust 3D-aware representation without multi-view or dynamic data, while ensuring consistent texture and geometry across different viewpoints. Although some baseline methods are capable of 3D-aware generation, the quality of the generated images still lags behind state-of-the-art 2D generation approaches, which excel in producing high-quality, detailed images. To address this severe limitation, we propose a novel feed-forward pipeline based on pixel-aligned Gaussian Splatting, coined as F3D-Gaus, which can produce more realistic and reliable 3D renderings from monocular inputs. In addition, we introduce a self-supervised cycle-consistent constraint to enforce cross-view consistency in the learned 3D representation. This training strategy naturally allows aggregation of multiple aligned Gaussian primitives and significantly alleviates the interpolation limitations inherent in single-view pixel-aligned Gaussian Splatting. Furthermore, we incorporate video model priors to perform geometry-aware refinement, enhancing the generation of fine details in wide-viewpoint scenarios and improving the model’s capability to capture intricate 3D textures. Extensive experiments demonstrate that our approach not only achieves high-quality, multi-view consistent 3D-aware generation from monocular datasets, but also significantly improves training and inference efficiency. Project Page: https://w-ted.github.io/publications/F3D-Gaus
@@ 1 Introduction
3D-aware generation from a single image has wide applications in virtual reality (VR), augmented reality (AR), gaming, and autonomous driving. Traditional methods for 3D-aware generation from a single image typically use voxels [9], point clouds [59] or meshes [47, 13] as 3D representations. However, in recent years, with the rise of NeRF [26] and 3D Gaussian Splatting (3DGS) [23] representations, methods based on NeRF [3, 39] and 3DGS [42, 44] have gained popularity due to their advantages in capturing fine details and handling complex lighting conditions.
@@
Figure 1: Illustration of our motivation for cycle-consistent self-supervised training. For monocular datasets: (a) supervision is naturally available for the canonical view. (b) For novel views, where supervision is absent, we use the rendered novel-view image as input to obtain its 3D representation. This 3D representation is then re-rendered from the canonical view, where supervision is available. Red arrows indicate feed-forward 3D representation prediction from a monocular image, while blue arrows represent the rendering processes from 3D representations at different specific viewpoints.
In general, most recent 3D-aware generation methods rely on multi-view image supervision [42, 44, 19, 65, 48, 35, 32], fine-tuning on video data [12, 62] , or optimization based on 2D/3D diffusion priors [43, 60]. However, compared to the availability of multi-view or high-resolution video data, monocular datasets are far more abundant and easier to collect. This is particularly advantageous because monocular data can be gathered from a wide range of sources, such as single-frame images from the web or consumer-grade cameras, without requiring specialized equipment or controlled capture conditions. Our approach aims to explore the possibility of performing 3D-aware generation tasks using only monocular datasets, without the need for multi-view images or video data during training. This monocular setting also presents the potential for scaling up to larger datasets, which can lead to better generalization across a wider variety of scenes and objects. On the other hand, inference-time optimization methods [12, 43, 60] for 3D-aware generation often suffer from low efficiency and are impractical for real-world applications, especially in time-sensitive applications or environments with limited computational capacity. Therefore, to enhance accessibility and practicality, we propose a feed-forward framework based on pixel-aligned Gaussian Splatting that aims to directly infer 3D-aware representations in one single step, using only monocular datasets for training.
However, since monocular datasets only provide supervision for the frontal view, obtaining constraints for novel viewpoints is a key challenge. Previous methods trained on monocular datasets typically employ discriminators [39], semantic/geometry constraints [30], or iterative in-painting techniques [54] to generate plausible novel view images. We observe that while pixel-aligned Gaussian Splatting has advantages over previous approaches in terms of rendering realism, it struggles when the viewpoint changes significantly. Specifically, regions occluded in the original view are not easily interpolated with monocular Gaussian primitives to render plausible color and geometry in the new view. To address this issue, we introduced a self-supervised cycle constraint during training. This strategy ensures that the 3D representations from novel viewpoints are aligned with and complementary to those from the original viewpoint. By enforcing this cycle constraint, the model is naturally able to extrapolate across views by aggregating multiple representations to enhance the 3D capability during inference. Additionally, to correct any inconsistencies in geometry or textures caused by large viewpoint shifts, we incorporate a video in-painting model for refinement, resulting in a more reliable final output.
Throughout the development and evaluation of our approach, our 3D rendering results exhibit a remarkable level of realism, representing a substantial improvement over current challenging benchmarks. We also validate the robustness of our approach on a diverse set of datasets, including several single-category object-centric datasets and the more varied and extensive ImageNet [10] dataset, confirming its generalizability and practical applicability. In addition, our solution surpasses baseline methods in computational efficiency, significantly reducing both training and inference times without compromising performance.
@@
Our work provides a holistic solution for realistic and efficient 3D-aware generation for monocular datasets, marking a significant advancement in 3D content generation. The summarized contribution of our paper is threefold:
@@
  • We pioneer 3D-aware generation using generalizable feed-forward Gaussian Splatting representation, achieving significant efficiency and favorable rendering quality on monocular datasets.
  • We significantly advance the capability of pixel-aligned Gaussian Splatting representations by designing a self-supervised cycle training strategy specifically tailored for monocular datasets.
@@
  • We further mitigate the artifacts of 3D-aware representations caused by large viewpoint shifts by introducing geometry-aware video priors.
@@ 2 Related Works
Figure 2: Illustration of our overall framework. Given a single RGB image $I_{0}$ and depth map $D_{0}$, our model directly feeds them forward to output the pixel-aligned Gaussian Splatting representation $GS_{0}$, which can be used for novel view synthesis. After obtaining the 3DGS representation, we render the image $\tilde{I_{1}}$ and depth maps $\tilde{D_{1}}$ for the novel view, and then output its corresponding 3DGS $GS_{1}$. These two 3DGS representations are subsequently aggregated to produce the images for supervision. This novel self-supervised training strategy enforces cycle-consistent 3D representation learning across different views, allowing the generalized 3DGS representations to reinforce each other, thereby collaboratively enhancing the overall 3D representation capability.
@@ 2.1 Novel view synthesis for single input
With multi-view dataset. Novel view synthesis from a single image can be achieved through various methods. One straightforward strategy is to train or fine-tune models using multi-view or video data with camera information, which data naturally provides strong multi-view cues for 3D-aware generation. For instance, multi-view data is used to train models from scratch [19, 32, 67, 42, 44], or fine-tune pre-trained models for image generation [48, 35, 12, 18] or video generation [15, 52, 25, 45, 62] to achieve camera motion control. Another category of methods involves scene-specific optimization by leveraging the prior knowledge of 2D/3D diffusion models [43, 28, 17, 60]. However, scene-specific approaches are not generalizable, as they are tailored to individual objects or scenes.
With only monocular dataset. The approach most similar to ours trains a generalized model solely on monocular datasets and supports feed-forward predictions [3, 38, 54, 31, 30]. 3DGP [38] pioneered 3D-aware generation on ImageNet by training a model on 1,000 classes using a 3D-aware GAN for RGB-D prediction. It introduces a discriminator to distinguish real and generated images and employs a depth-adapter to refine unreliable depth predictions. Similarly, G3DR [30] uses a generalized tri-plane representation and semantic/perceptual losses for novel view rendering. Our method operates in a similar setting, but combines pixel-aligned Gaussian Splatting with a cycle-consistent training strategy, leading to more realistic renderings against the tri-plane approaches. In contrast, IVID [54] generates RGB-D in-painting pairs from ImageNet using a 2D diffusion model with iterative in-painting. However, it struggles to generalize to novel object categories, as it relies on matching input classes to ImageNet, inevitably limiting its adaptability to new categories.
@@
Recently, 3D Gaussian Splatting (3DGS) [23] stands as a leading representation for novel view synthesis, enabling real-time rendering with state-of-the-art visual quality. 3DGS models the 3D scene as a collection of learnable 3D Gaussian primitives with 3D coordinates, opacity, anisotropic covariance, and color features. Subsequent works applies 3DGS to downstream tasks such as scene understanding [68, 64, 2, 34, 70], scene editing [57, 46, 6, 50, 66], and surface reconstruction [24, 51, 5, 53, 20, 55, 63]. However, 3DGS was originally designed for scene-specific novel view synthesis. Recent works introduced generalizable models for predicting 3DGS from single or multiple images [42, 44, 41, 4, 8, 7, 40]. For example, Splatter-Image [42] pioneered pixel-aligned 3DGS prediction from RGB images, enabling fast feed-forward training and inference. LGM [44] combines pre-trained multi-view image generation models with an image-to-3DGS model, achieving object-level 3DGS inference from a single image. Flash3D [41] integrates a depth estimation model for scene-level 3DGS prediction from a single image. While these methods allow feed-forward inference from a single image to 3DGS, they still rely on multi-view data for training or use multi-view image generation models as priors. In contrast, our approach focuses on generalized 3DGS prediction without requiring multi-view images for training.
@@ 3 The Proposed Framework: F3D-Gaus
As shown in Fig. 2, we use a U-Net-based generalized 3DGS model as the backbone. Given an input image $I_{0}$ and its depth map $D_{0}$, the model predicts a pixel-aligned 3DGS representation, which can render images and depth maps from arbitrary viewpoints [63]. The details of the generalized 3DGS model are provided in Sec. 3.1. As shown in Fig. 1, the rendered outputs can be supervised in the canonical view but the monocular dataset lacks supervision for novel views. We propose a cycle-consistent self-supervised training strategy in Sec. 3.2 to address this. Additionally, our geometry-guided refinement method introduced in Sec. 3.3 further identifies and refines artifacts.
@@ 3.1 Preliminary
3D Gaussian Splatting. Each Gaussian primitive is characterized by its 3D coordinates $\bm{\mu}$, color features $\bm{c}$, opacity, scale matrix $S$, and rotation matrix $R$. With these attributes, the Gaussians are defined by the covariance matrix $\Sigma=RSS^{T}R^{T}$ centered at point $\bm{\mu}$:
$G(\bm{x})=\exp^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\Sigma^{-1}(\bm{x}-\bm{\mu})}.$ (1)
The covariance matrix is projected onto the 2D plane following [71], allowing us to compute the projected Gaussian and apply alpha-blending to obtain the final color on the image plane:
@@
Pixel-aligned 3DGS. Splatter-Image [42] introduces a framework for predicting pixel-aligned 3DGS from RGB images. Given an input image $I_{0}$, the model predicts Gaussian attribute maps, including 3D coordinates, opacity, color features, scale, and rotation. For an input of size $H\times W$, the output attribute maps have the same spatial size, corresponding to $n=H\times W$ Gaussian primitives. The scale and rotation define the covariance matrix. Together with 3D coordinates, opacity, and color features, the 3DGS can render images and depth maps from arbitrary viewpoints.
As shown in Fig. 2, we use a similar U-Net-based generalized 3DGS model as the backbone. The final 3D coordinates of the 3DGS are computed by adding the input depth $D_{0}$ to the predicted offset.
3.2 Cycle-consistent Self-supervised Strategy
The lack of novel view supervision motivates us to propose a cycle-consistent self-supervised strategy in this section. It aims to ensure the multi-view consistency of the predicted 3DGS. It has two key components: complementary aggregation and cycle supervision. The details are given below.
@@ 3.2 Cycle-consistent Self-supervised Strategy
$\mathcal{L}_{\text{recon}}=\lVert\hat{I_{0}}-I_{0}\rVert_{1}+\lVert\hat{D_{0}}-D_{0}\rVert_{1}$ (3)
@@
Figure 3: Illustration of complementary aggregation in the proposed cycle-consistency self-supervised strategy. We guide complementary aggregation by leveraging the differences between the alpha maps of the two 3DGS from different viewpoints.
@@
The cycle loss $\mathcal{L}_{\text{cycle}}$ in Equ. 7 looks the same as the reconstruction loss in Equ. 3, but they correspond to sampled canonical and novel views, respectively. The predictions of $\hat{I}_{0}$ and $\hat{D}_{0}$ differ in each case. Due to our complementary aggregation design, the gradients from $\mathcal{L}_{\text{cycle}}$ backpropagate through both $GS_{0}$ and $GS_{1}$. We block gradient propagating to $GS_{0}$ via $\tilde{I}_{1}$ and $\tilde{D}_{1}$, as we found that allowing it hinders the learning process.
@@
Note that novel loss $\mathcal{L}_{\text{novel}}$ defined in Equ. 9 is not part of our cycle-consistent training strategy, but rather a component of our baseline method. Even in the baseline method mentioned in the ablation study in Sec. 4.3, the novel loss is also applied (to $\tilde{I_{1}}$ and $\tilde{D_{1}}$ at that time). In addition to the above losses, we apply a regularization loss to enforce smoothness in the rendered depth.
@@
Inference-time aggregation. During training, we use a cycle-consistent training strategy by aggregating GS representations from two views and aligning them via cycle supervision. This naturally extends to multi-view aggregation during inference, enhancing representational capacity. At inference time, given inputs $I_{0}$ and $D_{0}$, we first obtain $GS_{0}$. We then render eight images from $GS_{0}$ along a spiral trajectory. Next, we run another inference step to obtain eight distinct $GS_{1}$ representations, which are concatenated with $GS_{0}$ to form the final GS representation.

### v3

# F3D-Gaus: Feed-forward 3D-aware Generation on ImageNet with Cycle-Aggregative Gaussian Splatting

@@ Abstract
This paper tackles the problem of generalizable 3D-aware generation from monocular datasets, e.g., ImageNet [9]. The key challenge of this task is learning a robust 3D-aware representation without multi-view or dynamic data, while ensuring consistent texture and geometry across different viewpoints. Although some baseline methods are capable of 3D-aware generation, the quality of the generated images still lags behind state-of-the-art 2D generation approaches, which excel in producing high-quality, detailed images. To address this severe limitation, we propose a novel feed-forward pipeline based on pixel-aligned Gaussian Splatting, coined as F3D-Gaus, which can produce more realistic and reliable 3D renderings from monocular inputs. In addition, we introduce a self-supervised cycle-aggregative constraint to enforce cross-view consistency in the learned 3D representation. This training strategy naturally allows aggregation of multiple aligned Gaussian primitives and significantly alleviates the interpolation limitations inherent in single-view pixel-aligned Gaussian Splatting. Furthermore, we incorporate video model priors to perform geometry-aware refinement, enhancing the generation of fine details in wide-viewpoint scenarios and improving the model’s capability to capture intricate 3D textures. Extensive experiments demonstrate that our approach not only achieves high-quality, multi-view consistent 3D-aware generation from monocular datasets, but also significantly improves training and inference efficiency. Project Page: https://w-ted.github.io/publications/F3D-Gaus
@@ 1 Introduction
Generating 3D-aware content from a single image has numerous applications in augmented reality and gaming, enhancing immersive experiences. However, achieving 3D awareness typically requires substantial supervision. Recent 3D-aware generation methods rely on optimization based on 2D/3D diffusion priors [39, 57], multi-view image supervision [38, 40, 16, 62, 45, 34, 32], or fine-tuning on video data [10, 59]. Among these, inference-time optimization approaches [10, 39, 57] often suffer from inefficiency, making them impractical for real-world applications, particularly in time-sensitive scenarios or resource-constrained environments. Consequently, there is a growing demand for feed-forward models capable of directly inferring 3D structure from single images.
@@
Figure 1: Illustration of our motivation for cycle self-supervised training. For monocular datasets: (a) supervision is naturally available for the canonical view. (b) For novel views, where supervision is absent, we use the rendered novel-view image as input to obtain its 3D representation. This 3D representation is then re-rendered from the canonical view, where supervision is available. Red arrows indicate feed-forward 3D representation prediction from a monocular image, while blue arrows represent the rendering processes from 3D representations at different specific viewpoints.
While most successful attempts [16, 38, 32, 59] rely on multi-view or video data, such datasets are often difficult and costly to obtain. In contrast, monocular data sets, consisting of single-frame images, are significantly more abundant and easier to collect. They can be sourced from the web or captured with consumer-grade cameras without requiring specialized equipment or controlled environments. Moreover, the scalability of monocular datasets enables training on large-scale data, which can enhance generalization across diverse scenes and object categories. These considerations motivate us to explore a key research question: Can we design a feed-forward 3D-aware generation framework that is trained purely on monocular image datasets?
This problem has long been central to computer vision, closely related to 3D representation learning. Traditional methods for 3D-aware generation from a single image typically employ explicit 3D representations, such as voxels [8], point clouds [56], or meshes [44, 11, 51]. However, these approaches suffer from limitations in appearance modeling and often struggle with generalization beyond specific object categories. More recently, radiance field-based representations have revolutionized 3D-aware generation due to their ability to model both geometry and appearance in a unified manner. The advent of NeRF [26] has led to significant progress in learning 3D structures from monocular datasets [2, 35, 30]. However, the implicit nature of NeRF, where the entire 3D asset is represented by triplane-based feature encoding and a shared MLP decoder, limits rendering quality. A promising alternative is the recently emerging 3D Gaussian Splatting (3DGS) [22], which represents radiance fields using discrete Gaussian primitives. This formulation improves rendering efficiency and inspires a novel feedforward framework: pixel-aligned Gaussian Splatting, designed to infer 3D-aware representations in a single forward pass using only monocular training data.
Despite its advantages, training on monocular datasets presents a fundamental challenge: since supervision is available only for frontal views, generating plausible novel viewpoints is non-trivial. Previous monocular-based methods address this issue using discriminators [35], semantic/geometry constraints [30], or iterative in-painting techniques [52]. While pixel-aligned GS offers advantages in rendering realism, it struggles with significant viewpoint changes. The occluded regions in the original view are difficult to interpolate, leading to incomplete geometry and artifacts in novel views.
To tackle this challenge, we introduce a cycle-aggregative strategy. It ensures that 3D representations from novel viewpoints are both aligned with and complementary to those from the original viewpoint. By enforcing this cycle consistency through aggregation, the model naturally learns to extrapolate across views by fusing multiple representations, enhancing 3D coherence during inference. Additionally, we incorporate a video in-painting model [66] guided by geometry information to correct inconsistencies in geometry and texture caused by large viewpoint shifts, resulting in more reliable and visually consistent outputs.
Through extensive evaluation, our approach achieves state-of-the-art realism in 3D rendering, surpassing existing benchmarks. We validate its robustness across diverse datasets to demonstrate strong generalization. Furthermore, our method significantly improves computational efficiency, reducing both training and inference times while maintaining high-quality outputs.
@@
In summary, our work presents a comprehensive solution for realistic and efficient 3D-aware generation from monocular datasets, marking a significant advancement in 3D content generation. Our contributions are threefold:
@@
  • We pioneer 3D-aware generation from single image using a generalizable, feed-forward Gaussian Splatting representation, achieving both high efficiency and superior rendering quality on monocular datasets.
  • We introduce a self-supervised cycle-aggregative training strategy to enhance the capability of pixel-aligned Gaussian Splatting, enabling more robust novel view synthesis from monocular images.
@@
  • We mitigate artifacts caused by large viewpoint shifts by integrating geometry-aware video priors, further refining 3D-aware representations.
@@ 2 Related Works
Figure 2: Illustration of our overall framework. Given a single RGB image $I_{0}$ and depth map $D_{0}$, our model directly feeds them forward to output the pixel-aligned Gaussian Splatting representation $GS_{0}$, which can be used for novel view synthesis. After obtaining the 3DGS representation, we render the image $\tilde{I_{1}}$ and depth maps $\tilde{D_{1}}$ for the novel view, and then output its corresponding 3DGS $GS_{1}$. These two 3DGS representations are subsequently aggregated to produce the images for supervision. This novel self-supervised training strategy enforces cycle-aggregative 3D representation learning across different views, allowing the generalized 3DGS representations to reinforce each other, thereby collaboratively enhancing the overall 3D representation capability.
@@ 2.1 Novel view synthesis for single input
With multi-view dataset. Novel view synthesis from a single image can be achieved through various methods. One straightforward strategy is to train or fine-tune models using multi-view or video data with camera information, which data naturally provides strong multi-view cues for 3D-aware generation. Some methods train models from scratch using multi-view data [16, 32, 38, 40, 23, 64], while others fine-tune pre-trained models for image [45, 34, 10, 15] or video generation [12, 49, 25, 42, 59] to enable camera motion control. Another category of methods optimizes scene-specific representations using 2D/3D diffusion priors [39, 28, 14, 57]. However, these approaches lack generalization ability, as they are tailored to individual objects or scenes.
With only monocular dataset. The most similar approach to ours trains a generalized model solely on monocular datasets for feed-forward predictions [2, 35, 52, 31, 30]. 3DGP [35] pioneered 3D-aware generation on ImageNet using a 3D-aware GAN for RGB-D prediction, with a discriminator for realism and a depth-adapter for geometry refinement. G3DR [30] employs a generalized tri-plane representation with semantic and perceptual losses for novel view rendering. Our method operates in a similar setting, but combines pixel-aligned Gaussian Splatting with a cycle-aggregative training, achieving more realistic renderings than tri-plane approaches. IVID [52] uses 2D diffusion for iterative RGB-D in-painting, but struggles with novel object categories, as it relies on ImageNet class labels. Real3D [18] uses a self-supervised strategy for 3D asset generation. However, our approach differs in both setting and technical design: We introduce a generalizable feed-forward GS for monocular datasets, while Real3D relies on LRM [16] and combines monocular and multi-view data. We additionally propose a complementary aggregation design for pixel-aligned GS and integrate video priors to enhance the learning process.
@@
Recently, 3D Gaussian Splatting (3DGS) [22] stands as a leading representation for novel view synthesis, enabling real-time rendering with state-of-the-art visual quality. It models the scene as learnable 3D Gaussian primitives with 3D coordinates, opacity, anisotropic covariance, and color features. Recent works extend 3DGS to tasks like scene understanding [1, 61, 33, 65, 67], editing [55, 43, 5, 47, 63], and surface reconstruction [24, 48, 4, 50, 17, 53, 60]. However, 3DGS was originally designed for scene-specific novel view synthesis. To generalize 3DGS prediction from single or multiple images, recent works [38, 40, 37, 3, 7, 6, 36] have emerged. Splatter-Image [38] pioneered pixel-aligned 3DGS prediction from RGB images, enabling fast feed-forward training and inference. LGM [40] combines pre-trained multi-view image generation models with an image-to-3DGS model for object-level 3DGS inference from a single image. Flash3D [37] integrates depth estimation for scene-level 3DGS prediction from a single image. While these methods allow feed-forward inference from a single image to 3DGS, they still rely on multi-view data for training or use multi-view image generation models as priors. In contrast, our approach focuses on generalized 3DGS prediction without requiring multi-view images for training.
@@ 3 The Proposed Framework: F3D-Gaus
As shown in Fig. 2, we use a U-Net-based generalized 3DGS model as the backbone. Given an input image $I_{0}$ and its depth map $D_{0}$, the model predicts a pixel-aligned 3DGS representation, which can render images and depth maps from arbitrary viewpoints. The details of the generalized 3DGS model are provided in Sec. 3.1. As shown in Fig. 1, the rendered outputs can be supervised in the canonical view but the monocular dataset lacks supervision for novel views. We propose a cycle-aggregative self-supervised training strategy in Sec. 3.2 to address this. Additionally, our geometry-guided refinement method introduced in Sec. 3.3 further identifies and refines artifacts.
@@ 3.1 Preliminary
3D Gaussian Splatting. Each Gaussian primitive is characterized by its 3D coordinates $\bm{\mu}$, color features $\bm{c}$, opacity, scale matrix $S$, and rotation matrix $R$. With these attributes, the Gaussians are defined by the covariance matrix $\Sigma=RSS^{T}R^{T}$ centered at point $\bm{\mu}$ following $G(\bm{x})=\exp^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\Sigma^{-1}(\bm{x}-\bm{\mu})}$. The covariance matrix is projected onto the 2D plane following [68], allowing us to compute the projected Gaussian and apply alpha-blending to obtain the final color on the image plane:
@@
Pixel-aligned 3DGS. Splatter-Image [38] introduces a framework for predicting pixel-aligned 3DGS from images. Given an input image $I_{0}$, it predicts Gaussian attribute maps, including 3D coordinates, opacity, color features, scale, and rotation, matching the same input spatial size with $n=H\times W$ Gaussian primitives. Scale and rotation define the covariance matrix. Together with other attributes, images and depth maps can be rendered from arbitrary viewpoints. We use a similar U-Net model as the backbone, while 3DGS’s coordinates are computed by adding the input depth $D_{0}$ to the predicted offset.
3.2 Cycle-aggregative Self-supervised Strategy
The lack of novel view supervision motivates us to propose a cycle-aggregative self-supervised strategy in this section. It aims to ensure the multi-view consistency of the predicted 3DGS. It has two key components: complementary aggregation and cycle supervision. The details are given below.
@@ 3.2 Cycle-aggregative Self-supervised Strategy
$\mathcal{L}_{\text{recon}}=\lVert\hat{I_{0}}-I_{0}\rVert_{1}+\lVert\hat{D_{0}}-D_{0}\rVert_{1}.$ (2)
@@
Figure 3: Illustration of the proposed cycle-aggregative self-supervised strategy. We guide complementary aggregation by leveraging the differences between the alpha maps of the two 3DGS from different viewpoints.
@@
The cycle loss $\mathcal{L}_{\text{cycle}}$ in Equ. 6 looks the same as the reconstruction loss in Equ. 2, but they correspond to sampled canonical and novel views, respectively. The predictions of $\hat{I}_{0}$ and $\hat{D}_{0}$ differ in each case. Due to our complementary aggregation design, the gradients from $\mathcal{L}_{\text{cycle}}$ backpropagate through both $GS_{0}$ and $GS_{1}$. We block gradient propagating to $GS_{0}$ via $\tilde{I}_{1}$ and $\tilde{D}_{1}$, as we found that allowing it hinders the learning process.
@@
Note that novel loss $\mathcal{L}_{\text{novel}}$ defined in Equ. 8 is not part of our cycle-aggregative training strategy, but rather a component of our baseline method. Even in the baseline method mentioned in the ablation study in Sec. 4.3, the novel loss is also applied (to $\tilde{I_{1}}$ and $\tilde{D_{1}}$ at that time). In addition to the above losses, we apply a total-variation regularization loss to enforce smoothness in the rendered depth.
@@
Inference-time aggregation. During training, we use a cycle-aggregative training strategy by aggregating GS representations from two views and aligning them via cycle supervision. This naturally extends to two-view aggregation during inference, enhancing representational capacity. At inference time, given inputs $I_{0}$ and $D_{0}$, we first obtain $GS_{0}$. We then render an image from $GS_{0}$ from the novel view. Next, we run another inference step to obtain novel-view representation $GS_{1}$, which are concatenated with $GS_{0}$ to form the final GS representation.